# LLM Watermarking Using Mixtures and Statistical-to-Computational Gaps

### v1

@@ Abstract.
Department of Mathematics, UC Irvine
1. Introduction
Large Language Models (LLMs) have emerged as a powerful technology for generating human-like text [3, 18]. On one side, an LLM performs well if it produces text that closely resembles human writing. On the other side, malicious use of high-performance LLMs also bring undesirable consequences such as the spread of misinformation [17], misuse in education [18, 13], and data pollution [15, 16].
@@ 1. Introduction
Before describing watermarking in more details, we recall the concept of tokenization. In a nutshell, a word consists in small pieces of “sub-words” known as tokens. A LLM outputs each token sequentially by computing a probability distribution over a fixed set of possible tokens (vocabulary) and sampling the next token from it. The distribution of the next token varies from token to token as it depends on the previous tokens sampled, while the vocabulary remains fixed during the process of text generation.
@@
  • Undetectability: Any polynomial-time algorithm based solely on the text generated by the LLM fails to detect any change in the probability distributions used to generate the tokens.
@@
1.1. Related Work
Several watermarking schemes were proposed [2, 14, 22] for the closed setting without any formal guarantee. Perhaps, the first watermarking scheme with provable guarantees is from [9], where the authors proposed to split the vocabulary into a green list and red list. The probabilities corresponding to the tokens in the green list are slightly increased while the ones in the red list are slightly decrease. The watermarking can be detected by checking the frequency of tokens in the green list versus tokens in the red list. Therefore, the downside of this approach is that the “undetectability” requirement is not fulfilled.
Another line of work [6, 12, 1, 10] is dedicated to the following idea: Let $u$ be a random variable distributed uniformly over the interval $[0,1]$ and $p=(p(1),\ldots,p(d))$ be a probability distribution over a vocabulary of size $d$. We can sample the next token according to $p$ by sampling from $u$ first and then observing that for any $k\in\{1,\ldots,d\}$
@@ 1.1. Related Work
To the best of our knowledge, the only result for the open source setting is from [5]. The authors proposed to perturb each logit in the softmax rule (see equation 2.1) by a vector sampled from a multivariate Gaussian distribution and exploits the correlation between the text and such vector. The authors provided some theoretical evidence (partially rigorous) for completeness and unremovability under strong assumptions on the text. [^1] For example, [5] assumes that the token distribution behaves as an uniform distribution over a certain subset and also that the adversary makes changes respecting some normalization of the soft-max function which are hard to verify.
1.2. Main Contributions
Our main contributions are new connections between watermarking and robust statistics to derive more efficient watermarking schemes. Our first main result is Theorem 3.2, where we propose an elementary watermarking scheme for the closed setting satisfying all the requirements (undetectability, completeness and soundness) under mild assumptions on the distribution of the text. We also argue that the assumptions are necessary in some sense.
In a nutshell, our watermarking scheme proceeds as follows: In the first step it randomly constructs partitions of the vocabulary set into green and red tokens that change at each time a new token is sampled. Similarly to [9], the probabilities of the green tokens are shifted upwards, and the ones for the red tokens, downwards. However, we make some key changes to this shifting scheme, which allows it to achieve undetectability.
Our second main result Theorem 4.3 lies in the realm of the open source setting. By leveraging novel connections to the theory of statistical-to-computational gaps in robust statistics, we proposed a watermarking scheme that is both “sound” and “complete”, along with mathematically rigorous guarantees. This watermark is also “unremovable”: any algorithm that attempts to remove it must (indirectly) solve a computationally hard problem – the sparse mean estimation under Huber’s contamination model (see [8] for a comprehensive introduction).
@@ 1.3. Roadmap
2. The Hypothesis Testing Formulation
Let $T$ be our vocabulary of $d$ tokens, which we can identify with $[d]\coloneqq\{1,\ldots,d\}$ without loss of generality. A text is a sequence of random variables $x_{i}$ taking values in $[d]$. At each step, the LLM computes logits $L=(\ell(1),\ldots,\ell(d))$ and samples the next token $x\in[d]$ according to the softmax rule
(2.1) $p(i)\coloneqq\mathbb{P}\{x=i\}=\operatorname{softmax}(\ell(1),\ldots,\ell(d))\coloneqq\frac{e^{\ell(i)}}{e^{\ell(1)}+\cdots+e^{\ell(d)}}.$
@@ 2. The Hypothesis Testing Formulation
3. Watermarking Scheme – Closed Setting
@@
Let $N$ be the length of the text to be generated [^2] Of course, the exact value of $N$ is unknown but we can always work with the maximum number of tokens allowed by the LLM. and $d$ be the size of the token vocabulary. Without loss of generality we assume that $d$ is even. [^3] In the case that $d$ is odd, we simply add a spurious token Next, let $r$ be a Rademacher random variable (±$\pm 1$ with equal probabilities) and consider the alternating signs vector $\Delta\coloneqq(r,-r,r,-r,\ldots,r)$ of length $d$. The secret key consists on i.i.d. copies $\Delta_{1},\ldots,\Delta_{N}$ of the random vector $\Delta$, one for each token in the text $x_{1},\ldots,x_{N}$ to be watermarked.
For simplicity of notation, we drop the index $j$ while describing the $j$-th step of the sampling algorithm. At each new step, the LLM computes the probability distribution of the next token $x$, namely $p=(p(1),\ldots,p(d))$. Now the sampling algorithm watermarks this distribution by increasing the probabilities $p(i)$ for the tokens $i$ where $\Delta(i)>0$ and decreasing them where $\Delta(i)<0$.
To do that, we first assume that $p(1)\geq\cdots\geq p(d)$. Define for each $i\in[d]$:
(3.1) $\varepsilon_{i}=\begin{cases}p(i),\quad\text{$i$ is even}\\ p(i+1),\quad\text{$i$ is odd},\end{cases}$
so that these numbers are pairwise equal: $\varepsilon_{1}=\varepsilon_{2}$, $\varepsilon_{3}=\varepsilon_{4}$, …, $\varepsilon_{d-1}=\varepsilon_{d}$. Next, we compute the watermarked distribution by setting for every $i\in[d]$:
(3.2) $q(i)=p(i)+\varepsilon_{i}\Delta(i).$
In the general case, we first compute a non-increasing rearrangement of $p$, perturb it as before, and rearrange back.
Now let’s check that this watermarking scheme is valid (i.e. $q$ is actually a probability distribution) and undetectable.
Proposition 3.1 (Validity and undetectability).
For any probability distribution $p$ over $[d]$, the watermarked distribution $q$ computed in (3.2) is indeed a probability distribution over $[d]$, and the watermarking satisfies the undetectability requirement.
@@
Sine $\Delta$ alternates signs and $\varepsilon_{1},\ldots,\varepsilon_{d}$ are pairwise equal, we have
$\sum_{i=1}^{d}\varepsilon_{i}\Delta(i)=\Delta(1)\Big{(}\underbrace{(\varepsilon_{1}-\varepsilon_{2})}_{=0}+\underbrace{(\varepsilon_{3}-\varepsilon_{4})}_{=0}+\ldots+\underbrace{(\varepsilon_{d-1}-\varepsilon_{d})}_{=0}\Big{)}=0.$
It follows that
$\sum_{i=1}^{d}q(i)=\sum_{i=1}^{d}p(i)+\sum_{i=1}^{d}\varepsilon_{i}\Delta(i)=1.$
So, to check that $q$ defines a probability distribution, it remains to show that it has nonnegative entries. But this follows from our construction: $q(i)\geq p(i)-p(i)=0$ for even $i$, and $q(i)\geq p(i)-p(i+1)\geq 0$ for odd $i$.
Now let’s check undetectability. For every $i\in[d]$, the probability that $x=i$ under $q$ is
$\begin{split}&\mathbb{P}_{q}\{x=i\}=(p(i)+\varepsilon_{i})\mathbb{P}\{\Delta(i)=1\}+(p(i)-\varepsilon_{i})\mathbb{P}\{\Delta(i)=0\}\\ &=\frac{1}{2}(p(i)+\varepsilon_{i})+\frac{1}{2}(p(i)-\varepsilon_{i})=p(i).\end{split}$
@@
3.1. Watermarking Detection – Closed setting
We now describe how our detection algorithm works. The core idea is that whenever the perturbation $\Delta(x)$ is positive, it increases the chance of the token $x$ to appear in the text. Thus, it is more likely to observe the tokens in the text that correspond to the positive entries of $\Delta$.
So, our algorithm just counts the fraction of tokens $x_{j}$ in the text that have positive $\Delta_{j}(x_{j})$, and tests if it significantly exceeds $1/2$:
@@ 3.1. Watermarking Detection – Closed setting
The text $x_{1},\ldots,x_{N}$. The secret key: $\Delta_{1},\ldots,\Delta_{N}$. Tolerance $\delta$.
True or False. $Z\leftarrow\frac{1}{N}\sum_{j=1}^{N}{\mathds{1}}_{\{\Delta_{j}(x_{j})=1\}}$.
if $Z\geq 1/2(1+\sqrt{3\log(1/\delta)/N})$ then return: True
end ifreturn: False
Now, we state that the watermarking we just described is undetectable, sound, and complete:
Theorem 3.2 (Watermarking in Closed Setting).
The watermarking scheme described above satisfies the following properties:
  1. (1)
    It is undetectable.
Moreover, for any $\delta\in(0,1)$, if $N\geq 3\log(1/\delta)$, then:
  1. (2)
    It is sound with probability at least $1-\delta$.
  2. (3)
    Let $p^{*}_{j}$ denote the probability of the most likely token in the vocabulary at step $j=1,\ldots,N$. If, for some $\gamma>1$,
    (3.3) $1+\frac{1}{N}\sum_{i=1}^{N}(1-p^{*}_{j})\geq\gamma\left(1+\sqrt{\frac{3\log(1/\delta)}{N}}\right),$
    then the watermarking scheme is complete with probability at least $1-e^{-c_{\gamma}N}$, where $c_{\gamma}\coloneqq(1-\gamma)/8$.
The result follows by taking $t=\gamma/2$ and using (4.3).
@@
Proof.
We already proved undetectability in Proposition 3.1.
Soundness: Suppose we are under the null hypothesis $H_{0}$ that the text $x_{1},\ldots,x_{N}$ is independent of the secret key $\Delta_{1},\ldots,\Delta_{N}$. Then the test $Z$ defined in Algorithm 1 has binomial distribution $(N,1/2)$. By Chernoff’s inequality, for every $t>0$ we have
$\mathbb{P}\{Z\geq 1/2(1+t)\}\leq e^{-t^{2}N/3},$
which is at most $\delta$ for $t=\sqrt{3\log(1/\delta)/N}$.
Completeness: Assume that the text is watermarked. Thus, at each fixed step, a random token $x$ is picked from the distribution $q$ defined in (3.2). Without loss of generality, assume that $p(1)\geq p(2)\geq\cdots\geq p(d)$. Then,
$\mathbb{P}\{\Delta(x)=1\}$ $={\mathbb{E}}{\mathds{1}}_{\{\Delta(x)=1\}}={\mathbb{E}}\sum_{i=1}^{d}{\mathds{1}}_{\{\Delta(i)=1\}}{\mathds{1}}_{\{x=i\}}$
$={\mathbb{E}}\Bigg{[}\sum_{i=1}^{d}{\mathds{1}}_{\{\Delta(i)=1\}}\,{\mathbb{E}}\big{[}{\mathds{1}}_{\{x=i\}}\,|\,\Delta\big{]}\Bigg{]}$
@@
$={\mathbb{E}}\sum_{i=1}^{d}{\mathds{1}}_{\{\Delta(i)=1\}}\,\big{(}p(i)+\varepsilon(i)\big{)}$
(by (3.2), since only the terms with $\Delta(i)=1$ contribute to the sum)
$=\sum_{i=1}^{d}\mathbb{P}\{\Delta(i)=1\}\,\big{(}p(i)+\varepsilon(i)\big{)}=\frac{1}{2}\sum_{i=1}^{d}\big{(}p(i)+\varepsilon(i)\big{)}$
(since each $\Delta(i)$ has Rademacher distribution)
$=\frac{1}{2}\sum_{i=1}^{d}p(i)+\frac{1}{2}\sum_{i=1}^{d}\varepsilon(i)$
$=\frac{1}{2}+\frac{1}{2}\Big{(}p(2)+\underbrace{p(2)}_{\geq p(3)}+p(4)+\underbrace{p(4)}_{\geq p(5)}+\cdots+p(d)\Big{)}$
(by construction (3.1) and the monotonicity assumption)
$\geq\frac{1}{2}+\frac{1}{2}(1-p(1)).$
@@
$Z_{k}\coloneqq Z_{k-1}+\mathds{1}_{\{\Delta_{k}(x_{k})=1\}}-{\mathbb{E}}\mathds{1}_{\{\Delta_{k}(x_{k})=1|x_{1},\ldots,x_{k-1}\}},$
is a martingale satisfying that $|Z_{(k)}-Z_{(k-1)}|\leq 1$ almost surely. Thus, by combining the Azuma-Hoeffding’s inequality
$\mathbb{P}\left\{Z\geq\frac{1}{2}\bigg{(}1+\frac{1}{N}\sum_{i=1}^{N}(1-p^{*}_{j})\bigg{)}-t\right\}\geq 1-e^{-t^{2}N/2},$
with the assumption (3.3), the result follows for $t=(\gamma-1)/2$. ∎
4. Watermarking Scheme - Open Source Setting
Unfortunately, in the open source setting, there is no guarantee that an adversary cannot destroy our watermarking scheme. So we propose a completely different approach in this setting.

### v2

@@ Abstract
1 Introduction
Large Language Models (LLMs) have emerged as a powerful technology for generating human-like text [3, 20]. On one side, an LLM performs well if it produces text that closely resembles human writing. On the other side, the use of high-performance LLMs also bring undesirable consequences such as the spread of misinformation [19], misuse in education [20, 15], and data pollution [17, 18].
@@ 1 Introduction
Before describing watermarking in more details, we recall the concept of tokenization. In a nutshell, a word consists in small pieces of “sub-words” known as tokens. A LLM outputs each token sequentially by computing a probability distribution over a fixed set of possible tokens (dictionary) and sampling the next token from it. The distribution of the next token varies from token to token as it depends on the previous tokens sampled, while the dictionary remains fixed during the process of text generation.
@@
  • Statistical Undetectability: Any algorithm based solely on the text generated by the LLM fails to detect any change in the probability distributions used to generate the tokens  [^1] It should not be confused with the notion of cryptographic undetectability as in [7]..
@@
1.1 Related Work
Several watermarking schemes were proposed [2, 16, 24] for the closed setting without any formal guarantee. Perhaps, the first watermarking scheme with provable guarantees is from [11], where the authors proposed to split the dictionary into a green list and red list. The probabilities corresponding to the tokens in the green list are slightly increased while the ones in the red list are slightly decrease. The watermarking can be detected by checking the frequency of tokens in the green list versus tokens in the red list. Therefore, the downside of this approach is that the “undetectability” requirement is not fulfilled.
Another line of work [7, 14, 1, 12] is dedicated to the following idea: Let $u$ be a random variable distributed uniformly over the interval $[0,1]$ and $p=(p(1),\ldots,p(d))$ be a probability distribution over a dictionary of size $d$. We can sample the next token according to $p$ by sampling from $u$ first and then observing that for any $k\in\{1,\ldots,d\}$
@@ 1.1 Related Work
To the best of our knowledge, the only result for the open source setting is from [6]. The authors proposed to perturb each logit in the softmax rule (see equation 2.1) by a vector sampled from a multivariate Gaussian distribution (see [4] for a similar idea) and exploits the correlation between the text and such vector. The authors provided some theoretical evidence (partially rigorous) for completeness and unremovability under strong assumptions on the text.  [^2] For example, [6] assumes that the token distribution behaves as an uniform distribution over a certain subset and also that the adversary makes changes respecting some normalization of the soft-max function which are hard to verify.
1.2 Main Contributions
Our main contributions are new connections between watermarking and robust statistics to derive more efficient watermarking schemes. Our first main result follows from Proposition 3.3, Proposition 3.4 and Theorem 3.5, where we propose an elementary watermarking scheme for the closed setting satisfying all the requirements (undetectability, completeness and soundness) under mild assumptions on the distribution of the text. We also argue that the assumptions are necessary in some sense.
In a nutshell, our watermarking scheme proceeds as follows: In the first step it randomly constructs partitions of the dictionary set into green and red tokens that change at each time a new token is sampled. Similarly to [11], the probabilities of the green tokens are shifted upwards, and the ones for the red tokens, downwards. However, we make some key changes to this shifting scheme, which allows it to achieve undetectability and obtain explicit non-asymptotic guarantees that are easier to interpret than the more involved guarantees available in the literature.
Our second main result follows from Proposition 4.2, Proposition 4.3 and Theorem 4.4 and lies in the realm of the open source setting. By leveraging novel connections to the theory of statistical-to-computational gaps in robust statistics, we proposed a watermarking scheme that is both “sound” and “complete”, along with mathematically rigorous guarantees. This watermark is also “unremovable”: any algorithm that attempts to remove it must (indirectly) solve a computationally hard problem – the sparse mean estimation under Huber’s contamination model (see [10] for a comprehensive introduction).
@@ 1.3 Roadmap
2 The Hypothesis Testing Formulation
Let $T$ be our dictionary of $d$ tokens, which we can identify with $[d]\coloneqq\{1,\ldots,d\}$ without loss of generality. A text is a sequence of random variables $x_{i}$ taking values in $[d]$. At each step, the LLM computes logits $L=(\ell(1),\ldots,\ell(d))$ and samples the next token $x\in[d]$ according to the softmax rule
$p(i)\coloneqq\mathbb{P}\{x=i\}=\operatorname{softmax}(\ell(1),\ldots,\ell(d))\coloneqq\frac{e^{\ell(i)}}{e^{\ell(1)}+\cdots+e^{\ell(d)}}.$ (2.1)
@@ 2 The Hypothesis Testing Formulation
3 Closed Setting
@@
3.1 The watermarking scheme
Denote the dictionary (the set of all possible tokens) by $[d]=\{1,\ldots,d\}$ and assume that $d$ is even – otherwise, add a spurious token. To generate a text consisting of $N$ tokens, do the following for each step $j=1,\ldots,N$.
Randomly and independently paint the dictionary into two colors $1$ and $-1$ with exactly $d/2$ tokens receiving each color. Such a coloring is encoded by the random vector $\Delta_{j}\in\{-1,1\}^{d}$. Also, generate an independent Rademacher random variable $r_{j}$, whose purpose is to flip a coloring. The secret key consists of the colorings $r_{1}\Delta_{1},\ldots,r_{N}\Delta_{N}$.
Now we will describe the watermark at each given step $j=1,\ldots,N$, and to ease the notation we will drop the subscript $j$.
The LLM computes the probability distribution of the next token:
$p=(p(1),\ldots,p(d)).$
We will now watermark this distribution by increasing the probabilities $p(i)$ for the tokens $i$ where $r\Delta(i)=1$ and decreasing them where $r\Delta(i)=-1$.
To do that, let $I$ and $I^{c}$ denote the subsets of tokens colored $1$ and $-1$, respectively, i.e.
$I\coloneqq\{i\in[d]:\;\Delta(i)=1\}.$
Now we match the tokens of opposite colors according to their rank: the most likely token colored $1$ is matched with the most likely token colored $-1$, the second most likely with the second most likely, and so on. Then, if $r=1$, we increase the probability of the first element of each pair and decrease the probability of the second element. If $r=-1$, we do the opposite.
Formally, let $w(i)$ denote the $i$-th largest probability of a token with color $1$, i.e. the largest element of the set $\{p(i)\}_{i\in I}$. Similarly, let $\overline{w}(i)$ be the $i$-th largest element of the set $\{p(i)\}_{i\in I^{c}}$. Since each token $i\in[d]$ is ranked, there exists $i^{\prime}\in[d/2]$ such that $w(i^{\prime})=p(i)$ if $i\in I$ or $\overline{w}(i^{\prime})=p(i)$, otherwise. Set $\varepsilon(i)$ to be the smaller probability in each pair of tokens:
$\varepsilon(i)\coloneqq\min\big{(}w(i^{\prime}),\overline{w}(i^{\prime})\big{)}.$ (3.1)
Next, we compute the watermarked distribution by setting for every $i\in[d]$:
$q(i)=p(i)+\varepsilon(i)r\Delta(i).$ (3.2)
Sample token $x\in[d]$ from the probability distribution $q=(q(1),\ldots,q(d))$.
Repeat this procedure for steps $j=1,\ldots,N$ to generate a text string $x_{1},\ldots,x_{N}$.
Remark 3.1 (Practical Considerations).
In practice, we can fix just one random coloring $\Delta$ of the dictionary rather than generate an independent random coloring $\Delta_{j}$ for every next token. Independent colorings help us establish theoretical guarantees, but they could be too conservative in practice.
3.2 Validity and Undetectability
Now let’s check that the watermarking scheme we described is valid (i.e. $q$ is actually a probability distribution) and undetectable.
Proposition 3.2 (Validity).
The watermarked distribution $q$ defined in (3.2) is indeed a probability distribution over $[d]$.
@@ 3.2 Validity and Undetectability
By design, every token $i\in I$ is matched to some $j\in I^{c}$ such that $\varepsilon(i)=\varepsilon(j)$. Since $\Delta(i)$ and $\Delta(j)$ have opposite signs, we have $\Delta(i)\varepsilon(i)=-\Delta(j)\varepsilon(j)=0$. Thus, when we sum all these terms, they cancel:
$\sum_{i=1}^{d}\varepsilon(i)\Delta(i)=\sum_{i\in I}\Delta(i)\varepsilon(i)+\sum_{j\in I^{c}}\Delta(j)\varepsilon(j)=0,$
and we get
$\sum_{i=1}^{d}q(i)=\sum_{i=1}^{d}p(i)+r\sum_{i=1}^{d}\varepsilon(i)\Delta(i)=1.$
So, to check that $q$ defines a probability distribution, it remains to show that it has nonnegative entries. But this follows from our construction:
$q(i)\geq p(i)-\varepsilon(i)=p(i)-\min\left(w(i^{\prime}),\overline{w}(i^{\prime})\right)\geq p(i)-p(i)=0.\qed$
Proposition 3.3 (Undetectability).
The watermarked distribution $q$ defined in (3.2) is indeed a probability distribution over $[d]$, and the watermarking satisfies the undetectability requirement.
Proof.
For every $i\in[d]$, the probability that $x=i$ under $q$ is
$\begin{split}&\mathbb{P}_{q}\{x=i\}=(p(i)+r\varepsilon(i))\mathbb{P}\{\Delta(i)=1\}+(p(i)-r\varepsilon(i))\mathbb{P}\{\Delta(i)=-1\}\\ &=(p(i)+r\varepsilon(i))\frac{1}{2}+(p(i)-r\varepsilon(i))\frac{1}{2}=p(i).\end{split}$
@@
3.3 Soundness and Completeness
We now describe how our detection algorithm works. The idea is that whenever $r\Delta(x)=1$ is positive, the chance of the token $x$ to appear in the text is increased, and when $r\Delta(x)=1$ the chance is decreased.
Thus, the watermarked text should contain more tokens $x$ for which $r\Delta(x)=1$ than those for which $r\Delta(x)=-1$. This is testable: just check if the empirical mean of the random variables $r_{j}\Delta_{j}(x_{j})$ significantly exceeds zero.
On the other hand, an non-watermarked text is independent of the key, so it should contain roughly the same number of tokens $x$ for which $r\Delta(x)=1$ as those for which $r\Delta(x)=-1$. Thus, a non-watermarked text has a negligible chance to pass the test above.
Let’s make this all precise.
@@ 3.3 Soundness and Completeness
The text $x_{1},\ldots,x_{N}$. The secret key: $r_{1}\Delta_{1},\ldots,r_{N}\Delta_{N}$. Tolerance $\delta$.
True or False. Compute $Z=\frac{1}{N}\sum_{j=1}^{N}r_{j}\Delta_{j}\big{(}x_{j}\big{)}$. If $Z\geq\sqrt{2\log(1/\delta)/N}$ return True, else return False.
Proposition 3.4 (Soundness).
The watermarking scheme described in Section 3.1 is sound probability at least $1-\delta$.
Proof.
Suppose we are under the null hypothesis $H_{0}$ that the text $x_{1},\ldots,x_{N}$ is independent of the secret key $r_{1}\Delta_{1},\ldots,r_{N}\Delta_{N}$. Then the test $Z$ defined in Algorithm 1 is the normalized sum of independent bounded random variables. By Hoeffding’s inequality, for every $t>0$ we have
$\mathbb{P}\{Z\geq t\}\leq e^{-Nt^{2}/2},$
which is at most $\delta$ for $t=\sqrt{2\log(1/\delta)/N}$. ∎
Theorem 3.5 (Completeness).
Let $\delta\in(0,1)$ and $N\geq 2\log(1/\delta)$. Let $p^{*}_{j}$ denote the probability of the most likely token in the dictionary at step $j=1,\ldots,N$. If, for some $\gamma>0$,
$\frac{1}{20N}\sum_{i=1}^{N}(1-p^{*}_{j})\geq\sqrt{\frac{2\log(1/\delta)}{N}}+\gamma,$ (3.3)
then the watermarking scheme described in Section 3.1 is complete with probability at least $1-e^{-N\gamma^{2}/2}$.
@@
Before we proceed to the proof of Theorem 3.5, we require one preliminary lemma.
Lemma 3.7 (Maxima of subsets).
Consider any numbers $p_{1}\geq\cdots\geq p_{d}\geq 0$. Let $I$ be a random subset of $[d]$ with cardinality $d/2$, uniformly sampled from all such subsets. Denote by $w_{i}$ the $i$-th largest element of $\{p_{i}\}_{i\in I}$, and by $\overline{w}_{i}$ the $i$-th largest element of $\{p_{i}\}_{i\in I^{c}}$. Then
$\frac{1}{20}\sum_{i=2}^{d}p_{i}\leq{\mathbb{E}}\sum_{i=1}^{d/2}\min(w_{i},\overline{w}_{i})\leq\sum_{i=2}^{d}p_{i}.$
We prove this lemma in Appendix A.
Proof of Theorem 3.5.
Assume that the text is watermarked. Thus, at each step, given the previous tokens, the random token $x$ is picked from the distribution $q$ defined in (3.2). Then
${\mathbb{E}}r\Delta(x)={\mathbb{E}}\sum_{i=1}^{d}r\Delta(i){\mathds{1}}_{\{x=i\}}={\mathbb{E}}\sum_{i=1}^{d}r\Delta(i){\mathbb{E}}\Big{[}{\mathds{1}}_{\{x=i\}}|\Delta\Big{]}.$
By (3.2), we have
@@
So, using that $r\Delta(i)$ has Rademacher distribution, we get
${\mathbb{E}}r\Delta(x)=\sum_{i=1}^{d}p(i)\underbrace{{\mathbb{E}}r\Delta(i)}_{=0}+{\mathbb{E}}\sum_{i=1}^{d}\varepsilon(i)\underbrace{r^{2}\Delta(i)^{2}}_{=1}={\mathbb{E}}\sum_{i=1}^{d}\varepsilon(i).$
Now, using the definition of $\varepsilon(i)$ in (3.1) and applying Lemma 3.7, we get
${\mathbb{E}}r\Delta(x)={\mathbb{E}}\sum_{i=1}^{d/2}\min\big{(}w(i),\overline{w}(i)\big{)}\geq\frac{1}{20}\big{(}1-p^{*}\big{)},$
where $p^{*}=\max(p(1),\ldots,p(d))$. repeating this argument at each step $j=1,\ldots,N$ conditionally on the previous steps, we get
${\mathbb{E}}[r_{j}\Delta_{j}(x_{j})|x_{1},\ldots,x_{j-1}]\geq\frac{1}{20}\big{(}1-p^{\ast}_{j}\big{)}.$ (3.4)
@@
$Z_{j}\coloneqq Z_{j-1}+r_{j}\Delta_{j}(x_{j})-{\mathbb{E}}[r_{j}\Delta_{j}(x_{j})|x_{1},\ldots,x_{j-1}],$
is a martingale satisfying that $|Z_{k}-Z_{k-1}|\leq 1$ almost surely. Thus, by combining the Azuma-Hoeffding’s inequality and (3.4) we obtain
$\mathbb{P}\left\{Z\geq\frac{1}{20N}\sum_{i=1}^{N}(1-p^{*}_{j})-\gamma\right\}\geq 1-e^{-N\gamma^{2}/2}.$
Now it remains to use assumption (3.3). ∎
4 Open Setting
4.1 The watermarking scheme
Unfortunately, in the open source setting, there is no guarantee that an adversary cannot learn our watermarking scheme. So we propose a completely different approach in this setting.